two driverless trains race toward each other on the same track; one is traveling at a constant speed of 60 miles per hour and the other at 70 miles per hour. a fly starts at the front of one train and flies toward the second train. when he reaches it, he turns around and flies back to the first train, when he turns around, and so on. the fly maintains a constant speed of 110 miles per hour, losing no time in changing direction. if the trains are now 65 miles apart, how far will the fly fly before he is crushed when the trains collide?
Found 2 solutions by stanbon, Edwin McCravy:
Answer by stanbon(75887)

two driverless trains race toward each other on the same track; one is traveling at a constant speed of 60 miles per hour and the other at 70 miles per hour. a fly starts at the front of one train and flies toward the second train. when he reaches it, he turns around and flies back to the first train, when he turns around, and so on. the fly maintains a constant speed of 110 miles per hour, losing no time in changing direction. if the trains are now 65 miles apart, how far will the fly fly before he is crushed when the trains collide?
------------------------
The trains are converging at (70+60) = 130 mph
---
Time to go 65 miles = 65/130 = 1/2 hr.
----
The fly will fly (1/2 hr)(110 miles/hr) = 55 miles before being crushed.
=================
Cheers,
Stan H.
=================

Answer by Edwin McCravy(20064)

The trains are approaching each other at a rate of
60+70 or 130 mph. [That's how fast the original 
65 mile distance between them is shrinking.] Since 
time = distance/rate, the trains will collide in 
65/130 hours which reduces to 1/2 hour.  During that 
half hour, the fly, traveling at 110 mph, will have 
flown half of 110 miles or 55 miles.

Edwin